# Input Convex Lipschitz RNN: A Fast and Robust Approach for Engineering Tasks

### v5

@@ Abstract
Computational efficiency and robustness are essential in process modeling, optimization, and control for real-world engineering applications. While neural network-based approaches have gained significant attention in recent years, conventional neural networks often fail to address these two critical aspects simultaneously or even independently. Inspired by natural physical systems and established literature, input convex architectures are known to enhance computational efficiency in optimization tasks, whereas Lipschitz-constrained architectures improve robustness. However, combining these properties within a single model requires careful review, as inappropriate methods for enforcing one property can undermine the other. To overcome this, we introduce a novel network architecture, termed Input Convex Lipschitz Recurrent Neural Networks (ICLRNNs). This architecture seamlessly integrates the benefits of convexity and Lipschitz continuity, enabling fast and robust neural network-based modeling and optimization. The ICLRNN outperforms existing recurrent units in both computational efficiency and robustness. Additionally, it has been successfully applied to practical engineering scenarios, such as modeling and control of chemical process and the modeling and real-world solar irradiance prediction for solar PV system planning at LHT Holdings in Singapore.
@@
Input Convex Neural Networks, Lipschitz Constrained Neural Networks, Chemical Processes, Model Predictive Control
@@ 1 Introduction
In real-world applications, computational efficiency (for neural network-based optimization tasks) plays a pivotal role due to the imperative need for real-time decision-making, especially in critical engineering processes such as chemical process operations (Wang et al., 2025). We define computational efficiency as the speed at which computational tasks are completed within a given execution timeframe (in actual deployment to industrial engineering processes, the model is trained offline, and the costs associated with the training process are not a primary concern). An effective approach to achieving this efficiency is through the transformation from non-convex to convex structures. Convexity is a ubiquitous characteristic observed in various physical systems such as the potential energy function (Goldstein et al., 2002; Taylor and Taylor, 2005), magnetic fields (Purcell, 1963; Griffiths, 2005), and general relativity (Hartle, 2003).
@@
Additionally, in real-world engineering applications, robustness plays a critical role due to the prevalent noise present in most real-world data, which significantly hampers neural network performance. Our definition of robustness aligns closely with the principles outlined by Tan and Wu (2024). Given the inherent noise in real-world process data, our goal is to improve neural networks by effectively learning from noisy training data for comprehensive end-to-end applications. Leveraging LNNs offers a potential solution to the robustness challenge. However, at this stage, the development of neural networks that address the above two challenges simultaneously remains an open question. It is important to note that the integration of input convexity and Lipschitz continuity in a single neural network presents a non-trivial challenge due to their interplay in the way that enforcing one property in the design of neural networks may compromise the other property.
@@
Motivated by the above considerations, in this work, we introduce an Input Convex Lipschitz Recurrent Neural Network (ICLRNN) that amalgamates the strengths of input convexity and Lipschitz continuity. This approach aims to achieve the concurrent resolution of computational efficiency and robustness for real-world engineering applications. Specifically, we validate the efficacy of ICLRNN across various engineering problems that include neural network-based modeling and control for chemical processes and solar irradiance forecasting for energy systems. Performance assessment hinges on modeling accuracy, computational efficiency (i.e., evaluated by MPC runtime), model complexity (i.e., evaluated by the number of floating point operations (FLOPs)111 FLOPs, or floating-point operations, refer to the number of floating-point operations (such as addition, subtraction, multiplication, division, etc.) required to call the model once (i.e., a a single forward pass through a model). It serves as a metric for evaluating and quantifying the complexity and computational demand of the model, where higher FLOPs indicate greater complexity.), and robustness against noise. Our objective is to achieve equilibrium across these dimensions as they hold equal importance in real-world engineering scenarios.
@@
In summary, this work proposes an ICLRNN architecture, substantiating its theoretical attributes as input convex and Lipschitz continuous, and successfully implements and illustrates that ICLRNN surpasses state-of-the-art recurrent units in various engineering tasks, including modeling and control of chemical processes and real-time solar irradiance forecasting on real-world solar PV systems. The rest of this article is structured as follows: Section 2 introduces the class of systems and the design of neural network-based MPC. Section 3 provides an overview of the current family of recurrent units. Section 4 presents the proposed ICLRNN architecture and implementation, and proves that the architecture is both input convex and Lipschitz-constrained. Section 5 uses a chemical process example to validate the performance of ICLRNN in modeling and control ordinary differential equation (ODE) systems. Section 6 applies ICLRNN to solar irradiance forecasting on real-world solar PV systems.
@@ 2.1 Notations
[…]
@@ 2.1.1 Class of Systems
We consider a category of systems that can be expressed through the ordinary differential equations (ODEs) of the following form:
@@
where $\bm{\mathbf{u}}\in\mathbb{R}^{m}$ denotes the control action and $\bm{\mathbf{x}}\in\mathbb{R}^{n}$ denotes the state vector. The function $F:X\times U\to\mathbb{R}^{n}$ is continuously differentiable, where $X\subset\mathbb{R}^{n}$ and $U\subset\mathbb{R}^{m}$ are connected and compact subsets that enclose an open neighborhood around the origin respectively. In this context, we maintain the assumption throughout this task that $F(0,0)=0$, ensuring that the origin $(\bm{\mathbf{x}},\bm{\mathbf{u}})=(0,0)$ serves as an equilibrium point. As it might be impractical to obtain first-principles models for intricate real-world systems, our objective is to construct a neural network to model the nonlinear system specified by Eq. (1), embed it into optimization problems, and maintain computational efficiency and robustness simultaneously.
@@ 2.1.2 MPC Formulation
The optimization problem of the MPC that incorporates a ICLRNN as its predictive model is expressed as follows:
[…]
$\bm{\mathbf{u}}(t)\in\mathbb{U},\ \forall t\in[t_{k},t_{k+N})$ (2c)
@@
where $S(\Delta)$ denotes the set of piecewise constant functions with period $\Delta$, $\widetilde{\bm{\mathbf{x}}}$ denotes the predicted state trajectory, and $N$ denotes the number of sampling periods in the prediction horizon. The objective function $\mathcal{L}$ outlined in Eq. (2a) integrates a cost function $J$ that depends on both the control actions $\bm{\mathbf{u}}$ and the system states $\bm{\mathbf{x}}$. The system dynamic function $F_{nn}(\widetilde{\bm{\mathbf{x}}}(t),\bm{\mathbf{u}}(t))$ expressed in Eq. (2b) is parameterized by an RNN (e.g., the proposed ICLRNN in this work). Eq. (2c) encapsulates the constraint function $\mathbb{U}$, delineating feasible control actions. Eq. (2d) establishes the initial condition $\widetilde{\bm{\mathbf{x}}}(t_{k})$ in Eq. (2b), referring to the state measurement at $t=t_{k}$.
@@ 4.1 Introduction to Input Convex Lipschitz RNN
We propose a novel Input Convex Lipschitz Recurrent Neural Networks, termed ICLRNN, that possesses both Lipschitz continuity and input convexity properties. Specifically, the output of ICLRNN is defined as follows:
@@
where $\{\bm{\mathbf{W}}^{(x)},\bm{\mathbf{W}}^{(y)},\bm{\mathbf{U}}^{(x)}\}$ are constrained to be non-negative with largest singular values small and bounded by 1, and all $g_{i}$ are constrained to be convex, non-decreasing, and Lipschitz continuous (e.g., ReLU). Similar to Chen et al. (2018), we expand the input as $\hat{\bm{\mathbf{x}}}_{t}=\begin{bmatrix}\bm{\mathbf{x}}^{\top}_{t},-\bm{\mathbf{x}}^{\top}_{t}\end{bmatrix}^{\top}\in\mathbb{R}^{2n_{x}}$. Noted that the ICLRNN maintains the neural architecture of the simple RNN in Eq. (3) while introducing constraints on the weights and activation functions to ensure input convexity and Lipschitz continuity.
A neural network’s output inherits input convexity and Lipschitz continuity only if every hidden state possesses these properties. Unlike Chen et al. (2018) and Erichson et al. (2020), the ICLRNN enforces these properties by constraining the weights and activation functions rather than adding supplementary variables to the RNN architecture. This approach reduces network complexity and minimizes computational demands. Moreover, models such as LRNN, ICRNN, UniCORNN, and Linear Recurrent Units involve high FLOPs and lack theoretical guarantees on robustness and computational efficiency for neural network-based optimization problems. Many engineering applications can achieve satisfactory modeling accuracy with relatively simple neural network structures. However, improving the robustness and computational efficiency of these models for neural network-based optimization and control remains an important challenge. These considerations motivated the design choice to retain the standard RNN architecture while imposing input convexity and Lipschitz constraints on the weights.
4.2 Implementation of ICLRNN
@@
. Next, we perform spectral normalization on the non-negative weights, reducing the spectral norm of individual weights to at most 1, i.e.,
@@
, where $\epsilon$ is a small constant to ensure that operator norm of normalized weight is strictly less than 1, e.g., $1\times 10^{-3}$. To approximate $\sigma_{max}(\bm{\mathbf{W}})$, we use the power iteration algorithm (see Algorithm 1), an iterative numerical method that computes the largest singular value by repeatedly multiplying the matrix by a vector and normalizing the result until convergence. The complete code is available in https://github.com/killingbear999/ICLRNN. While spectral normalization has been widely used to stabilize the training of generative adversarial networks (GANs) (Miyato et al., 2018), it is less commonly applied to RNNs.
@@
Input: $maxiter$: maximum number of iterations
Input: $eps$: convergence tolerance (stopping criterion)
@@
2:  Normalize $\bm{\mathbf{v}}$: $\bm{\mathbf{v}}\leftarrow\frac{\bm{\mathbf{v}}}{||\bm{\mathbf{v}}||}$.
@@
5:     Normalize $\bm{\mathbf{u}}$: $\bm{\mathbf{u}}\leftarrow\frac{\bm{\mathbf{u}}}{||\bm{\mathbf{u}}||}$.
@@
7:     Normalize $\bm{\mathbf{v}}_{new}$: $\bm{\mathbf{v}}_{new}\leftarrow\frac{\bm{\mathbf{v}}_{new}}{||\bm{\mathbf{v}}_{new}||}$.
8:     Check for convergence: if $||\bm{\mathbf{v}}_{new}-\bm{\mathbf{v}}||<eps$, break.
@@
11:  Compute the largest singular value: $\sigma_{max}\leftarrow||\bm{\mathbf{W}}\bm{\mathbf{v}}||$.
@@
Moreover, Serrurier et al. (2021) employs the Björck algorithm to ensure Lipschitz continuity by iteratively refining an initial estimate of an orthogonal matrix. This algorithm minimizes the error between the estimate and the true orthogonal matrix, gradually converging to a stable solution, and adjusts singular values to at most one. However, it is important to emphasize that the Björck algorithm should not be used in the ICLRNN. Unlike spectral normalization, which scales weights to ensure the spectral norm is bounded and inherently maintains non-negativity, the Björck algorithm introduces both positive and negative values during its orthogonalization process. This behavior is incompatible with the non-negativity constraint required in ICLRNN.
Additionally, regularization techniques such as dropout (Srivastava et al., 2014), batch normalization (Ioffe, 2015), and layer normalization (Ba, 2016) can help stabilize neural network training. However, these techniques introduce non-linear transformations, disrupting network convexity and thus should not be used in the training of ICNNs (e.g., the training of our ICLRNN).
At last, ICLRNN uses ReLU as the activation function for hidden states instead of the GroupSort function proposed in Anil et al. (2019). This decision aims to maintain convexity within the model architecture. Although GroupSort exhibits gradient norm preservation and higher expressive power (Tan and Wu, 2024), its non-convex nature contrasts with our primary objective.
@@
For ICNNs, including our proposed ICLRNN, we recommend applying MinMax scaling to the data prior to fitting it to the neural network. MinMax scaling is a normalization technique that transforms data to a specified range, usually between 0 and 1 (or another specified range). This technique linearly scales each feature $x$ in the dataset $\mathcal{X}$ individually based on its minimum and maximum values, i.e., $\forall x\in\mathcal{X}$:
$x\leftarrow\frac{x-\min(\mathcal{X})}{\max(\mathcal{X})-\min(\mathcal{X})}(b-a)+a$ (10)
where $\min(\mathcal{X})$ is the minimum value of the feature in the dataset, $\max(\mathcal{X})$ is the maximum value of the feature in the dataset, $a$ is the lower bound of the desired range (default is 0), and $b$ is the upper bound of the desired range (default is 1). This preprocessing step ensures that all feature values lie within the specified range, which can improve the stability and performance of the neural network.
4.3 Lipschitz Continuity of ICLRNN
We first list the definition of Lipschitz continuity and related lemmas for readers’ reference, as follows:

### v6

# Input Convex Lipschitz Recurrent Neural Networks for Robust and Efficient Process Modeling and Optimization

@@ Abstract
Computational efficiency and robustness are essential in process modeling, optimization, and control for real-world engineering applications. While neural network-based approaches have gained significant attention in recent years, conventional neural networks often fail to address these two critical aspects simultaneously or even independently. Inspired by natural physical systems and established literature, input convex architectures are known to enhance computational efficiency in optimization tasks, whereas Lipschitz-constrained architectures improve robustness. However, combining these properties within a single model requires careful review, as inappropriate methods for enforcing one property can undermine the other. To overcome this, we introduce a novel network architecture, termed Input Convex Lipschitz Recurrent Neural Networks (ICL-RNNs). This architecture seamlessly integrates the benefits of convexity and Lipschitz continuity, enabling fast and robust neural network-based modeling and optimization. The ICL-RNN outperforms existing recurrent units in both computational efficiency and robustness. Additionally, it has been successfully applied to practical engineering scenarios, such as chemical process modeling and the modeling and control of Organic Rankine Cycle-based waste heat recovery systems.
@@
Input Convex Neural Networks, Lipschitz Constrained Neural Networks, Chemical Processes, Model Predictive Control, Organic Rankine Cycle, Waste Heat Recovery
@@ 1 Introduction
In real-world applications, computational efficiency (for neural network-based optimization tasks) plays a pivotal role due to the imperative need for real-time decision-making, especially in critical engineering processes such as chemical process operations (Wang et al., 2025). We define computational efficiency as the speed at which computational tasks are completed within a given execution timeframe (in actual deployment to industrial engineering processes, the model is trained offline, and the costs associated with the training process are not a primary concern). An effective approach to achieving this efficiency is through the transformation from non-convex to convex structures. Convexity is a ubiquitous characteristic observed in various physical systems such as the potential energy function (Goldstein et al., 2002; Taylor, 2005), magnetic fields (Purcell, 1963; Griffiths, 2005), and general relativity (Hartle, 2003).
@@
Additionally, in real-world engineering applications, robustness plays a critical role due to the prevalent noise present in most real-world data, which significantly hampers neural network performance. Our definition of robustness aligns closely with the principles outlined by (Tan and Wu, 2024). Given the inherent noise in real-world process data, our goal is to improve neural networks by effectively learning from noisy training data for comprehensive end-to-end applications. Leveraging LNNs offers a potential solution to the robustness challenge. However, at this stage, the development of neural networks that address the above two challenges simultaneously remains an open question. It is important to note that the integration of input convexity and Lipschitz continuity in a single neural network presents a non-trivial challenge due to their interplay in the way that enforcing one property in the design of neural networks may compromise the other property.
@@
Motivated by the above considerations, in this work, we introduce an Input Convex Lipschitz Recurrent Neural Network (ICL-RNN) that amalgamates the strengths of input convexity and Lipschitz continuity. This approach aims to achieve the concurrent resolution of computational efficiency and robustness for real-world engineering applications. Specifically, we validate the efficacy of ICL-RNN across various engineering problems that include neural network-based modeling and control for chemical processes and energy systems. Performance assessment hinges on modeling accuracy, computational efficiency (i.e., evaluated by MPC runtime), model complexity (i.e., evaluated by the number of floating point operations (FLOPs)  [^1] FLOPs, or floating-point operations, refer to the number of floating-point operations (such as addition, subtraction, multiplication, division, etc.) required to call the model once (i.e., a a single forward pass through a model). It serves as a metric for evaluating and quantifying the complexity and computational demand of the model, where higher FLOPs indicate greater complexity.), and robustness against noise. Our objective is to achieve equilibrium across these dimensions as they hold equal importance in real-world engineering scenarios.
@@
In summary, this work proposes an ICL-RNN architecture, substantiating its theoretical attributes as input convex and Lipschitz continuous, and successfully implements and illustrates that ICL-RNN surpasses state-of-the-art recurrent units in various engineering tasks, including modeling and control of processes in continuous stirred tank reactor (CSTR) systems and Organic Rankine Cycle (ORC)-based waste heat recovery systems. The rest of this article is structured as follows: Sec. 2 introduces the class of systems and the design of neural network-based MPC. Sec. 3 provides an overview of the current family of recurrent units. Sec. 4 presents the proposed ICL-RNN architecture and implementation, and proves that the architecture is both input convex and Lipschitz-constrained. Sec. 5 uses an example of a chemical process in CSTR systems to validate the performance of ICL-RNN in modeling and control of an ordinary differential equation (ODE) system. Sec. 6 applies ICL-RNN to the modeling and control of an ORC-based waste heat recovery system.
@@ 2.1 Notations
Let $L_{f}$ represent the Lipschitz constant of a function $f$. $\circ$" denotes composition. $\|\cdot\|$ and $\|\cdot\|_{2}$ denote the spectral norm and Euclidean norm of a matrix, respectively. $\sigma(\cdot)$ denotes the singular value of a matrix.
@@ 2.1.1 Class of Systems
We consider a class of nonlinear systems that can be expressed by the ordinary differential equations (ODEs) of the following form:
@@
where $\bm{\mathbf{u}}\in\mathbb{R}^{m}$ denotes the control action and $\bm{\mathbf{x}}\in\mathbb{R}^{n}$ denotes the state vector. The function $F:\mathbb{X}\times\mathbb{U}\to\mathbb{R}^{n}$ is continuously differentiable, where $\mathbb{X}\subset\mathbb{R}^{n}$ and $\mathbb{U}\subset\mathbb{R}^{m}$ are connected and compact subsets that enclose an open neighborhood around the origin respectively. In this context, we maintain the assumption throughout this task that $F(0,0)=0$, ensuring that the origin $(\bm{\mathbf{x}},\bm{\mathbf{u}})=(0,0)$ serves as an equilibrium point. As it might be impractical to obtain first-principles models for intricate real-world systems, our objective is to construct a neural network to model the nonlinear system specified by Eq. (1), embed it into optimization problems, and maintain computational efficiency and robustness simultaneously.
@@ 2.1.2 MPC Formulation
The optimization problem of the MPC that incorporates an ICL-RNN as its predictive model is expressed as follows:
$\mathcal{J}=\min_{\bm{\mathbf{u}}\in S(\Delta)}\int_{t_{k}}^{t_{k+N}}L(\widetilde{\bm{\mathbf{x}}}(t),\bm{\mathbf{u}}(t))dt$ (2a)
$\text{s.t. }\dot{\widetilde{\bm{\mathbf{x}}}}(t)=F_{nn}(\widetilde{\bm{\mathbf{x}}}(t),\bm{\mathbf{u}}(t))$ (2b)
$\bm{\mathbf{x}}(t)\in\mathbb{X},\bm{\mathbf{u}}(t)\in\mathbb{U},\ \forall t\in[t_{k},t_{k+N})$ (2c)
@@
where $S(\Delta)$ denotes the set of piecewise constant functions with period $\Delta$, $\widetilde{\bm{\mathbf{x}}}$ denotes the predicted state trajectory, and $N$ denotes the number of sampling periods in the prediction horizon. The objective function $\mathcal{J}$ outlined in Eq. (2a) integrates a cost function $L$ that depends on both the control actions $\bm{\mathbf{u}}$ and the system states $\bm{\mathbf{x}}$. The system dynamic function $F_{nn}(\widetilde{\bm{\mathbf{x}}}(t),\bm{\mathbf{u}}(t))$ expressed in Eq. (2b) is parameterized by an RNN (e.g., the proposed ICL-RNN in this work). Eq. (2c) encapsulates the constraint function $\mathbb{U}$, delineating feasible control actions. Eq. (2d) establishes the initial condition $\widetilde{\bm{\mathbf{x}}}(t_{k})$ in Eq. (2b), referring to the state measurement at $t=t_{k}$.
@@ 4.1 Introduction to Input Convex Lipschitz RNN
We propose a novel Input Convex Lipschitz Recurrent Neural Networks, termed ICL-RNN, that possesses both Lipschitz continuity and input convexity properties. Specifically, the output of ICL-RNN is defined as follows:
@@
where $\{\bm{\mathbf{W}}^{(x)},\bm{\mathbf{W}}^{(y)},\bm{\mathbf{U}}^{(x)}\}$ are constrained to be non-negative with largest singular values small and bounded by 1, and all $g_{i}$ are constrained to be convex, non-decreasing, and Lipschitz continuous (e.g., ReLU). Similar to Chen et al. (2018), we expand the input as $\hat{\bm{\mathbf{x}}}_{t}=\begin{bmatrix}\bm{\mathbf{x}}^{\top}_{t},-\bm{\mathbf{x}}^{\top}_{t}\end{bmatrix}^{\top}\in\mathbb{R}^{2n_{x}}$. Noted that the ICL-RNN maintains the neural architecture of the simple RNN in Eq. (3) while introducing constraints on the weights and activation functions to ensure input convexity and Lipschitz continuity.
A neural network’s output inherits input convexity and Lipschitz continuity only if every hidden state possesses these properties. Unlike Chen et al. (2018) and Erichson et al. (2020), the ICL-RNN enforces these properties by constraining the weights and activation functions rather than adding supplementary variables to the RNN architecture. This approach reduces network complexity and minimizes computational demands. Moreover, models such as LRNN, ICRNN, UniCORNN, and Linear Recurrent Units involve high FLOPs and lack theoretical guarantees on robustness and computational efficiency for neural network-based optimization problems. Many engineering applications can achieve satisfactory modeling accuracy with relatively simple neural network structures. However, improving the robustness and computational efficiency of these models for neural network-based optimization and control remains an important challenge. These considerations motivated the design choice to retain the standard RNN architecture while imposing input convexity and Lipschitz constraints on the weights.
4.2 Implementation of ICL-RNN
@@
Next, we perform spectral normalization on the non-negative weights, reducing the spectral norm of individual weights to at most 1, i.e.,
@@
where $\epsilon$ is a small constant to ensure that operator norm of normalized weight is strictly less than 1, e.g., $1\times 10^{-3}$. To approximate $\sigma_{max}(\bm{\mathbf{W}})$, we use the power iteration algorithm (see Algorithm 1), an iterative numerical method that computes the largest singular value by repeatedly multiplying the matrix by a vector and normalizing the result until convergence. The complete code is available in https://github.com/killingbear999/ICLRNN. While spectral normalization has been widely used to stabilize the training of generative adversarial networks (GANs) (Miyato et al., 2018), it is less commonly applied to RNNs.
@@
Input: maxitermaxiter: maximum number of iterations
Input: epseps: convergence tolerance (stopping criterion)
@@
[…]
@@
[…]
@@
[…]
8:  Check for convergence: if $\bm{\mathbf{v}}_{new}-\bm{\mathbf{v}}||<eps$, break.
@@
[…]
@@
Moreover, Serrurier et al. (2021) employs the Björck algorithm to ensure Lipschitz continuity by iteratively refining an initial estimate of an orthogonal matrix. This algorithm minimizes the error between the estimate and the true orthogonal matrix, gradually converging to a stable solution, and adjusts singular values to at most one. However, it is important to emphasize that the Björck algorithm should not be used in the ICL-RNN. Unlike spectral normalization, which scales weights to ensure the spectral norm is bounded and inherently maintains non-negativity, the Björck algorithm introduces both positive and negative values during its orthogonalization process. This behavior is incompatible with the non-negativity constraint required in ICL-RNN.
Additionally, regularization techniques such as dropout (Srivastava et al., 2014), batch normalization (Ioffe and Szegedy, 2015), and layer normalization (Ba et al., 2016) can help stabilize neural network training. However, these techniques introduce non-linear transformations, disrupting network convexity and thus should not be used in the training of ICNNs (e.g., the training of our ICL-RNN).
At last, ICL-RNN uses ReLU as the activation function for hidden states instead of the GroupSort function proposed in Anil et al. (2019). This decision aims to maintain convexity within the model architecture. Although GroupSort exhibits gradient norm preservation and higher expressive power (Tan and Wu, 2024), its non-convex nature contrasts with our primary objective.
@@
For ICNNs, including our proposed ICL-RNN, we recommend applying MinMax scaling to the data prior to fitting it to the neural network. MinMax scaling is a normalization technique that transforms data to a specified range, usually between 0 and 1 (or another specified range). This technique linearly scales each feature $x$ in the dataset $\mathbb{X}$ individually based on its minimum and maximum values, i.e., $\forall x\in\mathbb{X}$:
$x\leftarrow\frac{x-\min(\mathbb{X})}{\max(\mathbb{X})-\min(\mathbb{X})}(b-a)+a$ (10)
where $\min(\mathbb{X})$ is the minimum value of the feature in the dataset, $\max(\mathbb{X})$ is the maximum value of the feature in the dataset, $a$ is the lower bound of the desired range (default is 0), and $b$ is the upper bound of the desired range (default is 1). This preprocessing step ensures that all feature values lie within the specified range, which can improve the stability and performance of the neural network.
4.3 Lipschitz Continuity of ICL-RNN
We first present the definition of Lipschitz continuity and related lemmas for readers’ reference, as follows: